OpenStudy (anonymous):

Choose the correct simplification of the expression a2 • a3. a a6 a5 a−1

OpenStudy (anonymous):

is it a6

OpenStudy (blurbendy):

a^5 a^2 * a^3 = a^5 multiplying exponents with the same base means you just add them.

OpenStudy (anonymous):

thanx

OpenStudy (blurbendy):

np

OpenStudy (jhannybean):

\[\large a^x \cdot a^x = a^{x+x} = (a^{x})^2\]

OpenStudy (anonymous):

so a5 is the answer

OpenStudy (blurbendy):

yup